OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment whose endpoints are given. A (2, 3) and B (6, 11)

OpenStudy (anonymous):

Midpoint= (4,7) x-coordinate= (2+6)/2=4 y-coordinate=(3+11)/2=7